OpenStudy (anonymous):

given any two functions f(x) and g(x), is it always true that (f ºg)(x)=(gºf)(x)?

OpenStudy (kc_kennylau):

Try f(x)=x+2 and g(x)=2x

OpenStudy (anonymous):

but one of the answer choices are yes no or cannot be determined

OpenStudy (amistre64):

kc gave you a way to determine it ....

OpenStudy (anonymous):

so it the answer yes?

OpenStudy (anonymous):

is*

OpenStudy (amistre64):

did you compose the function of the define f and g to compare them?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

and what are the compositions you came up with?

OpenStudy (anonymous):

2=x

OpenStudy (amistre64):

f(x)=x+2 and g(x)=2x fog = g+2 = 2x+2 gof = 2f = 2(x+2) = 2x+4 is fog always equal to gof ? is 2x+2, always equal to 2x+4 ?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

then the answers has to be no ...

OpenStudy (anonymous):

ok I get it thank you

OpenStudy (amistre64):

by default, function composition is not commutative